OpenStudy (anonymous):

solve the differential equation dy/dx = x^2-y^2

OpenStudy (amistre64):

hmm, can we let vx = y ?

OpenStudy (amistre64):

v'x + v = x^2 - (vx)^2 v'x + v = x^2(1 - v^2) prolly not ...

OpenStudy (anonymous):

this appeared in the section before substitution, so I feel like an idiot not knowing the answer.

OpenStudy (amistre64):

its not a linear setup, and its not seperable that i can tell

OpenStudy (anonymous):

the answer in the book says you can guarantee the existence of a solution and the uniqueness of a solution

OpenStudy (amistre64):

y' = x^2-y^2 y' + y^2 = x^2 subbing in z = y^2 z' = 2yy' doesnt help since there is no y term

OpenStudy (anonymous):

theorem 1 says if dy/dx and Dy(dy/dx) is continuous on some interval near a, it guarantees uniqueness. so here, it seems its continuous everywhere. but for a solution to exist, don't you have to explicitly solve for y(x)?

OpenStudy (amistre64):

you dont need to explicitly solve for it no

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%27%2By%5E2%3Dx%5E2 the solution is a mess for sure

OpenStudy (anonymous):

but the examples in the book always solve it to show existence. am I misunderstanding?

OpenStudy (amistre64):

and how does the book present the question?

OpenStudy (amistre64):

x^2 - y^2 is a continuous function, it has no bad inputs to restrict Fy = -2yy' is just as good ..... so the thrm holds

OpenStudy (anonymous):

determine if theorem 1 guarantees existence. if it does, determine if theorem 1 guarantees the solution is unique.

OpenStudy (amistre64):

so, by thrm 1 .... we see that f(x,y) is good, and Dy f(x,y) is good

OpenStudy (amistre64):

havent you previously proved the thrm?

OpenStudy (anonymous):

I see the existence part. But does it prove uniqueness?

OpenStudy (amistre64):

you can always make the rectangle "domain" small enough so that f(x,y) is unique given these conditions

OpenStudy (anonymous):

for dy/dx = y^1/3 the answer says it proves existence of y(0)=0 but not uniqueness. if you take f(x,y) = y^1/3 then df/dy = 1/(3y^2/3) but thats not continuous at y=0. so how does that work?

OpenStudy (amistre64):

since Dy is not continuous at x=0, then the thrm is not wrong is it?

OpenStudy (amistre64):

IF f(x,y) and Dy are continuous at a, then we can be assured that there exists a unique solution for some y(a)=b

OpenStudy (anonymous):

that what I'm confused about

OpenStudy (amistre64):

spose we have 2 solutions ...  they can be defined as existing at their common point, and they can be defined in a unique fashion at every other point. It is only at the intersection that they are simply not uniquely definable